# Mixed Precision Block-Jacobi Preconditioner: Algorithms, Performance Evaluation and Feature Analysis

### v2

@@ 4.3 Performance Evaluation
$Speedup=\frac{T_{uniform}}{T_{mix}}$ (4.4)
where $T_{uniform}$ denotes the CPU time of the uniform precision BJAC PCG algorithm and $T_{mix}$ denotes the CPU time of the mixed precision BJAC PCG algorithm, i.e., fMP-BJAC or aMP-BJAC PCG algorithm. For aMP-BJAC(hl) and aMP-BJAC(lh), the performance is depend on the adaptive threshold ${adp}_{tol}$. In the experiments in this sections, unless otherwise statement, $adp_{tol}=10$ for aMP-BJAC(hl), and $adp_{tol}={10}^{-5}$ for aMP-BJAC(lh).
@@
For the Diff3D problems, fp64 is chosen as the high precision and fp32 is chosen as the low precision. From Figure 1, we can see that different mixed precision preconditioner have different performance behavior, and the performance gains differ from diffusion coefficients. For all cases, fMP-BJAC and aMP-BJAC(hl) have achieved comparable speedup, that of 1.3$\times$ to 1.5$\times$, which outperform the speedup of aMP-BJAC(lh).
For the RHD3D problems, fp80 is chosen for the high precision and fp32 for the low precision. From Figure 2, we can conclude that the fMP-BJAC and aMP-BJAC(hl) can achieved speedup of more than 1.5$\times$ for RHD3D-1T and 1.8$\times$ for RHD3D-3T respectively, and fMP-BJAC slightly outperforms aMP-BJAC(hl) for the RHD3D-1T case. Again, the speedups of both fMP-BJAC and aMP-BJAC(hl) are better than that of aMP-BJAC(lh) for two cases. In addition, we also investigate the results of fMP-BJAC with fp64 as the low precision (fp64-fMP-BJAC PCG in Figure 2), its speedup does not outperform that of fp32-fMP-BJAC PCG for both cases.
From the above results, it can be concluded that the mixed precision preconditioners fMP-BJAC and aMP-BJAC(hl) achieve performance gains compared to the uniform precision BJAC. Meanwhile, two issues should be noticed. Firstly, the adaptive precision with two orders, i.e., aMP-BJAC(hl) and aMP-BJAC(lh), have quite performance differences, and the "high-to-low" strategy is significantly better than the "low-to-high" strategy. Secondly, the mixed precision preconditioners have different performance gains for different problems. The first issue concerns the influence of the mixed precision strategy on the convergence behavior, and the second issue involves the impact of the application features on the performance of the mixed precision preconditioner. We furtherly analyse and discuss these two issues in the following section.
@@ 4.4.1 Convergence Delays Phenomenon
From the results of Diff3D problems in Figure 3, it can be observed that the convergence behavior of mixed precision preconditioners is different due to diffusion coefficients. For Diff3D-Ani(1000) and Diff3D-Rand(1000) cases, the mixed precision BJAC have the same iteration numbers as the fp64-uniform-BJAC, which achieve the perfect performance. However, for other two problems, Diff3D-Const and Diff3D-Dis(1000), the iteration numbers of the mixed precision preconditioners increases slightly compared to the fp64-uniform-BJAC. Concretely, for Diff3D-Const, the three mixed precision preconditioners increase iteration numbers by 11%, which is from 148 to 165. For Diff3D-Dis(1000), the convergence behavior is different due to preconditioners, fp32 fMP-BJAC and aMP32 MP-BJAC(lh) increase the iteration numbers by 14% which is from 185 to 211, and the fp32 aMP-BJAC(hl) increases the iteration numbers by 11% which is from 185 to 205.
@@ 4.4.2 Improving Convergence via Adaptive Precision Preconditioners
For three test cases, include Diff3D-Const, Diff3D-Dis(1000) and RHD3D-1T, where the mixed precision BJAC preconditioner causes convergence delay issue as shown in last section, Table 2 gives the results of convergence behavior, i.e., changing trend of iteration numbers, by varying the threshold ${adp}_{tol}$ with typical values.
@@
Figure 5: RelResNorm curves of the aMP-BJAC(hl) and aMP-BJAC(lh) for solving the RHD3D-1T problem with three typical values of ${adp}_{tol}$. Left: the overall result. Right: the detailed result after the 500th iteration.
As shown in Figure 5, in the early and middle stages during the iterations, the RelResNorm curves of six preconditioners exactly overlap. When the RelResNorm is reduced to around $5\times{10}^{-10}$, only the RelResNorm of the aMP-BJAC(hl) with ${adp}_{tol}={10}^{-1}$ continue to reduce, however, for other values of ${adp}_{tol}$, the RelResNorm curves occur bifurcation and reach the convergence criterion with additional iterations, i.e., the convergence delay occur. In particular, for the aMP-BJAC(lh), even if ${adp}_{tol}$ is increased to $10$, at which only the beginning 1st, the 57th-60th, and the 62nd-102nd iterations use low precision, and the rest of the iterations use high precision, the convergence delay still occur.
@@ 4.5 Feature Analysis for Convergence Delays
Convergence delay leads to a challenge for achieving ideal speedup of mixed preconditioners. An interesting observation is that, the three problems occurring the convergence delay, include Diff-Const, Diff-Dis(1000), and RHD3D-1T, are more closer to a model problem compared to other three test problems, and should intuitively more likely to achieve the ideal speedup, whereas the numerical results are just the opposite. This is an interesting phenomenon need to be concerned in the future.
In this section, we investigate the correlation between the convergence delay behaviour and the features of the problems. As shown in Table 1, the four Diff3D test cases and the RHD3D-1T problem have the same matrix sparse pattern as well as bandwidth since all tests use the 3D 7-point scheme as discretization. For the RHD3D-3T test problem, using the same discretization mesh and scheme as other test problems, however, its matrix size is three times larger than other test problems since the coupling of three-temperatures involved on each mesh cell. In addition, for all test problems, the resulting matrices are belong to M-matrix[2].
@@ 4.5.1 Multiscale Feature
$\tau_{i}=\frac{\max\{|a_{ij}|\text{,}i\neq j\}}{\min\{|a_{ij}|\text{,}i\neq j\}}.$ (4.5)
@@
Table 3 gives the distribution of the multiscale strength for the six problems. The results illustrate that Diff3D-Const is a single-scale problem, while Diff3D-Dis(1000) and RHD3D-1T have relatively weak multiscale strength, and the other problems have strong multiscale. Associate with the results in Section 4.4 (see Figure 3 and Figure 4), it can be found that the convergence delay of mixed precision preconditioners are related to the multiscale strength. Concretely, the cases occurring convergence delay are exactly single-scale or weak-multiscale problems. On the contrary, the strong multiscale problems, include Diff3D-Ani(1000), Diff3D-Rand(1000), and RHD3D3T, do not occur convergence delay.
@@
[100,   1000)1001000[100,\;\;\;1000)[ 100 , 1000 ) 0.0% 100.0% 0.0% 0.0% 0.0% 0.0%
$[1000,\;\;+\infty)$ 100.0% 0.0% 0.0% 0.0% 0.0% 0.0%
@@
Figure 6 show the decline of relative residuals norm(RelResNorm) as well as the true RelResNorm for the uniform-BJAC and the fp32-fMP-BJAC with the six different $s$. Here, true RelResNorm uses residual computed via the residual definition formula $r=b-Ax$, instead of the residual deduced in PCG iterations. From Figure 6, it can be found that for problem Diff3D-Ani(1000) , the two RelResNorm curves of the uniform-BJAC and the fp32-fMP-BJAC overlap completely, which indicates that the fp32-fMP-BJAC and the uniform-BJAC converge with the same convergence speed. For the problems Diff3D-Ani(100), Diff3D-Ani(10), Diff3D-Ani(4), Diff3D-Ani(2) and Diff3D-Ani(1), the two RelResNorm curves of the uniform-BJAC and fp32-fMP-BJAC are overlapped in the early iterations stage, but the bifurcations occur when the RelResNorms fall to $10^{-16}$, $10^{-11}$, $10^{-10}$, $10^{-9}$ and $10^{-10}$, respectively, and then the RelResNorm curves of fp32-fMP-BJAC suddenly rise and then fall again, finally reach to the convergence criterion. Since the convergence threshold is $10^{-10}$, for the problems Diff3D-Ani(1000), Diff3D-Ani(100), Diff3D-Ani(10), Diff3D-Ani(4), the number of converged iterations of the fMP-BJAC is the same as the uniform-BJAC, which are 327, 321, 205 and 179, respectively. It indicates that there is no delay phenomenon. However, for the Diff3D-Ani(2) and the Diff3D-Ani(1), i.e., Diff3D-Const, the number of iterations for these two preconditioners are different, 159 and 148 for uniform-BJAC while 177 and 165 for the fp32-fMP-BJAC, respectively, which show the convergence delay.
Similarly, for two practical problems, RHD3D-1T and RHD3D-3T, the multiscale strength distributions are shown in Table 3. It shows that the multiscale strength of RHD3D-1T is weaker than that of RHD3D-3T. The RelResNorm curves of the uniform-BJAC, the fp64-fMP-BJAC, and fp32-fMP-BJAC are given in Figure 7, it can be found that for both RHD3D-1T and RHD3D-3T, the RelResNorm curves of the uniform-BJAC and the fp64-fMP-BJAC overlap, which indicates that the fp64-fMP-BJAC does not lead to convergence delay. For RHD3D-1T, the RelResNorm curves of the uniform-BJAC and the fp32-fMP-BJAC bifurcate when RelResNorm fall to $10^{-9}$, and the number of iterations are 628 and 824, respectively. Since the convergence threshold is $10^{-12}$, that is, the fp32-fMP-BJAC leads to the convergence delay. For the RHD3D-3T problem, however, the RelResNorm curves of the uniform-BJAC and the fp32-fMP-BJAC overlap, indicating that the fp32-fMP-BJAC does not cause convergence delay.
@@ 4.5.2 Diagonal Dominance Feature
Diagonal Dominance is another common feature in many applications. A typical scene that leading to diagonal dominance is time-dependent problems, e.g. RHD3D-1T and RHD3D-3T problems considered in this paper, as shown in (4.2). For time-dependent problems, the items of time discretization with a time step $dt$ will be add to the diagonal entries of matrix that lead to the diagonal dominance, and the strength of the diagonal dominance is closely related to the time step $dt$, the smaller $dt$, the stronger the diagonal dominance, and vice versa. Theoretical analyses [2] show that the convergence of the Jacobi method is relevant to the strength of the diagonal dominance of the matrix.
@@
We observe the trend of changing in convergence delay as $P_{diag}$ increasing, the result is given in Figure 8, where $iter_{o}=iter_{m}/iter_{u}$ denotes the ratio of the iteration numbers of the fp32-fMP-BJAC to that of the uniform BJAC preconditioners, $iter_{o}$ reflects the overhead of additional iterations introduced by mixed precision preconditioner. As shown in Figure 8, as the diagonal dominance strength increasing, $iter_{o}$ shows a fluctuating trend of declining. For example, the $iter_{o}$ is 1.36, 1.24, and 1.20 when $P_{diag}$ is 1, 2, and 8, respectively, suggesting that an increase in the diagonal dominance strength can improve the convergence delay of the fMP-BJAC preconditioner.
@@ 5 Conclusions
Acknowledgments
This work is supported by the National Natural Science Foundation of China (No.62032023 and No.12301476), the Science Challenge Project of China (No.TZZT20240303) and National Key Research and Development Program of China (No.2023YFB3001605).

### v3

@@ 4.3 Performance Evaluation
$Speedup=\frac{T_{uniform}}{T_{mix}}$ (7)
where $T_{uniform}$ denotes the CPU time of the uniform precision BJAC PCG algorithm and $T_{mix}$ denotes the CPU time of the mixed precision BJAC PCG algorithm, i.e., fMP-BJAC or aMP-BJAC PCG algorithm. For aMP-BJAC(hl) and aMP-BJAC(lh), the performance is depend on the adaptive threshold ${adp}_{tol}$. In the experiments in this section, unless otherwise statement, $adp_{tol}=10$ for aMP-BJAC(hl), and $adp_{tol}={10}^{-5}$ for aMP-BJAC(lh).
@@
For the Diff3D problems, fp64 is chosen as the high precision and fp32 is chosen as the low precision. From Figure 1, we can see that different mixed precision preconditioner have different performance behavior, and the performance gains differ from diffusion coefficients. For all cases, fMP-BJAC and aMP-BJAC(hl) have achieved comparable speedup, that of 1.3x to 1.5x, which outperform the speedup of aMP-BJAC(lh).
For the RHD3D problems, fp80 is chosen for the high precision and fp32 for the low precision. From Figure 2, we can conclude that the fMP-BJAC and aMP-BJAC(hl) can achieve speedup of more than 1.5$\times$ for RHD3D-1T and 1.8$\times$ for RHD3D-3T respectively, and fMP-BJAC slightly outperforms aMP-BJAC(hl) for the RHD3D-1T case. Again, the speedups of both fMP-BJAC and aMP-BJAC(hl) are better than that of aMP-BJAC(lh) for two cases. In addition, we also investigate the results of fMP-BJAC with fp64 as the low precision (fp64-fMP-BJAC PCG in Figure 2), its speedup does not outperform that of fp32-fMP-BJAC PCG for both cases.
From the above results, it can be concluded that the mixed precision preconditioners fMP-BJAC and aMP-BJAC(hl) achieve performance gains compared to the uniform precision BJAC. Meanwhile, two issues should be noticed. Firstly, the adaptive precision with two orders, i.e., aMP-BJAC(hl) and aMP-BJAC(lh), have quite performance differences, and the ”high-to-low” strategy is significantly better than the ”low-to-high” strategy. Secondly, the mixed precision preconditioners have different performance gains for different problems. The first issue concerns the influence of the mixed precision strategy on the convergence behavior, and the second issue involves the impact of the application features on the performance of the mixed precision preconditioner. We furtherly analyse and discuss these two issues in the following section.
@@ 4.4.1 Convergence Delays Phenomenon
From the results of Diff3D problems in Figure 3, it can be observed that the convergence behavior of mixed precision preconditioners varies with diffusion coefficients. For Diff3D-Ani(1000) and Diff3D-Rand(1000) cases, the mixed precision BJAC have the same iteration numbers as the fp64-uniform-BJAC, which achieve the perfect performance. However, for other two problems, Diff3D-Const and Diff3D-Dis(1000), the iteration numbers of the mixed precision preconditioners increases slightly compared to the fp64-uniform-BJAC. Concretely, for Diff3D-Const, the three mixed precision preconditioners increase iteration numbers by 11%, which is from 148 to 165. For Diff3D-Dis(1000), the convergence behavior is different due to preconditioners, fp32 fMP-BJAC and aMP32 MP-BJAC(lh) increase the iteration numbers by 14% which is from 185 to 211, and the fp32 aMP-BJAC(hl) increases the iteration numbers by 11% which is from 185 to 205.
@@ 4.4.2 Improving Convergence via Adaptive Precision Preconditioners
For three test cases, include Diff3D-Const, Diff3D-Dis(1000) and RHD3D-1T, where the mixed precision BJAC preconditioners cause convergence delays as shown in last section, Table 2 gives the results of convergence behavior, i.e., changing trend of iteration numbers, by varying the threshold ${adp}_{tol}$ with typical values.
@@
Figure 5: RelResNorm curves of the aMP-BJAC(hl) and aMP-BJAC(lh) for solving the RHD3D-1T problem with three typical values of ${adp}_{tol}$. Upper panel: the overall result. Lower panel: the detailed result after the 500th iteration.
As shown in Figure 5, in the early and middle stages during the iterations, the RelResNorm curves of six preconditioners exactly overlap. When the RelResNorm is reduced to around $5\times{10}^{-10}$, only the RelResNorm of the aMP-BJAC(hl) with ${adp}_{tol}={10}^{-1}$ continue to reduce, however, for other values of ${adp}_{tol}$, the RelResNorm curves occur bifurcation and reach the convergence criterion with additional iterations, i.e., the convergence delays occur. In particular, for the aMP-BJAC(lh), even if ${adp}_{tol}$ is increased to $10$, at which only the beginning 1st, the 57th-60th, and the 62nd-102nd iterations use low precision, and the rest of the iterations use high precision, the convergence delays still occur.
@@ 4.5 Feature Analysis for Convergence Delays
Convergence delay leads to a challenge for achieving ideal speedup of mixed preconditioners. An interesting observation is that, the three problems occurring the convergence delay, include Diff-Const, Diff-Dis(1000), and RHD3D-1T, are more closer to a model problem compared to other three test problems, and should intuitively more likely to achieve the ideal speedup, whereas the numerical results are just the opposite. This is an interesting phenomenon that should be concerned about in the future.
In this section, we investigate the correlation between the convergence delays behaviour and the problem features. As shown in Table 1, the four Diff3D test cases and the RHD3D-1T problem have the same matrix sparse pattern as well as bandwidth since all tests use the 3D 7-point scheme as discretization. For the RHD3D-3T test problem, using the same discretization mesh and scheme as other test problems, however, its matrix size is three times larger than other test problems since the coupling of three-temperatures involved on each mesh cell. In addition, for all test problems, the resulting matrices are belong to M-matrix[2].
@@ 4.5.1 Multiscale Feature
$\tau_{i}=\frac{\max\{|a_{ij}|\text{,}i\neq j\}}{\min\{|a_{ij}|\text{,}i\neq j\}},\text{ \; when \;}a_{ij}\neq 0.$ (8)
@@
Table 3 gives the distribution of the multiscale strength for the six problems. The results illustrate that Diff3D-Const is a single-scale problem, while Diff3D-Dis(1000) and RHD3D-1T have relatively weak multiscale strength, and the other problems have strong multiscale. Associate with the results in Section 4.4 (see Figure 3 and Figure 4), it can be found that the convergence delays of mixed precision preconditioners are related to the multiscale strength. Concretely, the cases occurring convergence delays are exactly single-scale or weak-multiscale problems. On the contrary, the strong multiscale problems, including Diff3D-Ani(1000), Diff3D-Rand(1000), and RHD3D3T, do not occur convergence delays.
@@
[100,   1000)1001000[100,\;\;\;1000)[ 100 , 1000 ) 0% 100% 0% 0% 0% 0%
$[1000,\;\;+\infty)$ 100% 0% 0% 0% 0% 0%
Figure 6 show the decline of relative residuals norm(RelResNorm) as well as the true RelResNorm for the uniform-BJAC and the fp32-fMP-BJAC with the six different $s$. Here, true RelResNorm uses residual computed via the residual definition formula $r=b-Ax$, instead of the residual deduced in PCG iterations.
@@
From Figure 6, it can be found that for problem Diff3D-Ani(1000) , the two RelResNorm curves of the uniform-BJAC and the fp32-fMP-BJAC overlap completely, which indicates that the fp32-fMP-BJAC and the uniform-BJAC converge with the same convergence speed. For the problems Diff3D-Ani(100), Diff3D-Ani(10), Diff3D-Ani(4), Diff3D-Ani(2) and Diff3D-Ani(1), the two RelResNorm curves of the uniform-BJAC and fp32-fMP-BJAC are overlapped in the early iterations stage, but the bifurcations occur when the RelResNorms fall to $10^{-16}$, $10^{-11}$, $10^{-10}$, $10^{-9}$ and $10^{-10}$, respectively, and then the RelResNorm curves of fp32-fMP-BJAC suddenly rise and then fall again, finally reach to the convergence criterion. Since the convergence threshold is $10^{-10}$, for the problems Diff3D-Ani(1000), Diff3D-Ani(100), Diff3D-Ani(10), Diff3D-Ani(4), the number of converged iterations of the fMP-BJAC is the same as the uniform-BJAC, which are 327, 321, 205 and 179, respectively. It indicates that there is no delay phenomenon. However, for the Diff3D-Ani(2) and the Diff3D-Ani(1), i.e., Diff3D-Const, the number of iterations for these two preconditioners are different, 159 and 148 for uniform-BJAC while 177 and 165 for the fp32-fMP-BJAC, respectively, which show the convergence delay.
Similarly, for two practical problems, RHD3D-1T and RHD3D-3T, the multiscale strength distributions are shown in Table 3. It shows that the multiscale strength of RHD3D-1T is weaker than that of RHD3D-3T. The RelResNorm curves of the uniform-BJAC PCG, the fp64-fMP-BJAC PCG, and fp32-fMP-BJAC PCG are given in Figure 7. It can be found that for both RHD3D-1T and RHD3D-3T, the RelResNorm curves of the uniform-BJAC and the fp64-fMP-BJAC overlap, which indicates that the fp64-fMP-BJAC does not lead to convergence delay. For RHD3D-1T, the RelResNorm curves of the uniform-BJAC and the fp32-fMP-BJAC bifurcate when RelResNorm fall to $10^{-9}$, and the number of iterations are 628 and 824, respectively. Since the convergence threshold is $10^{-12}$, that is, the fp32-fMP-BJAC leads to the convergence delay. For the RHD3D-3T problem, however, the RelResNorm curves of the uniform-BJAC and the fp32-fMP-BJAC overlap, indicating that the fp32-fMP-BJAC does not cause convergence delay.
@@ 4.5.2 Diagonal Dominance Feature
Diagonal Dominance is another common feature in many applications. A typical scene that leading to diagonal dominance is time-dependent problems, e.g. RHD3D-1T and RHD3D-3T problems considered in this paper, as shown in (5). For time-dependent problems, the items of time discretization with a time step $dt$ will be add to the diagonal entries of matrix that lead to the diagonal dominance, and the strength of the diagonal dominance is closely related to the time step $dt$, the smaller $dt$, the stronger the diagonal dominance, and vice versa. Theoretical analyses [2] show that the convergence of the Jacobi method is relevant to the strength of the diagonal dominance of the matrix.
@@
We observe the trend of changing in convergence delay as $P_{diag}$ increasing, the result is given in Figure 8, where $iter_{o}=iter_{m}/iter_{u}$ denotes the ratio of the iteration numbers of the fp32-fMP-BJAC to that of the uniform BJAC preconditioners, $iter_{o}$ reflects the overhead of additional iterations introduced by mixed precision preconditioner.
@@
As shown in Figure 8, as the diagonal dominance strength increasing, $iter_{o}$ shows a fluctuating trend of declining. For example, the $iter_{o}$ is 1.36, 1.24, and 1.20 when $P_{diag}$ is 1, 2, and 8, respectively, suggesting that an increase in the diagonal dominance strength can improve the convergence delay of the fMP-BJAC preconditioner.
@@ 5 Conclusions
[…]
Acknowledgements
This research was funded by National Key Research and Development Program of China (No.2023YFB3001605), Science Challenge Project of China (No.TZ2024009), National Natural Science Foundation of China (No.62032023 and No.12301476).
Conflict of Interest
On behalf of all authors, the corresponding author states that there is no confict of interest.